OpenStudy (anonymous):

solve for x. 2/x=6/2 choices: 2/3 1 1/2 6

OpenStudy (anonymous):

3

OpenStudy (texaschic101):

2/x = 6/2 cross multiply 6x = 4....now solve for x

OpenStudy (anonymous):

2/3?

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

ok i have a few things about triangles aswell.. which polygons are similar. picture coming

OpenStudy (isaiah.feynman):

Well, two shapes are similar if they have AT LEAST two same angles.